OpenStudy (anonymous):

Evaluate exactly the value of the integral from negative 1 to 0 of the product of the cube of the quantity 2 times x to the 4th power plus 8 times x and 4 times x to the 3rd power plus 4, dx. Your work must include the use of substitution and the antiderivative.

OpenStudy (anonymous):

https://gyazo.com/a94388763f348d36040ce04fad93ecc0

OpenStudy (anonymous):

@jim_thompson5910 u substitution with multiplication rule?

OpenStudy (anonymous):

If so how would I approach it using u substitution?

OpenStudy (zarkon):

\[\Large u=2x^4+8x\]

OpenStudy (anonymous):

@Zarkon du would be equal to 8x^3

OpenStudy (zarkon):

no

OpenStudy (anonymous):

oops wrote down +8 instead of 8x

OpenStudy (anonymous):

So it would be 8x^3+8

OpenStudy (anonymous):

So I need to multiply by 2?

OpenStudy (zarkon):

almost

OpenStudy (anonymous):

Then solve?

OpenStudy (zarkon):

\[\Large du=(8x^3+8)dx\]

OpenStudy (anonymous):

I usually forget to include that thanks

OpenStudy (zarkon):

divide by 2

OpenStudy (anonymous):

Multiply the part in the integral by 2

OpenStudy (anonymous):

No?

OpenStudy (zarkon):

\[\Large \frac{du}{2}=(4x^3+4)dx\] then substitute

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

\[2(4x^3 + 4) = du\]

OpenStudy (zarkon):

\[2(4x^3 + 4)dx = du\] so \[(4x^3 + 4)dx = \frac{du}{2}\]

OpenStudy (anonymous):

Why is dividing by 2 preferred?

OpenStudy (zarkon):

because \((4x^3+4)dx\) is in the problem

OpenStudy (zarkon):

you can now replace that part with du/2

OpenStudy (anonymous):

I now have \[\int\limits_{-1}^{0}\frac{ u^3du}{ 2}\]

OpenStudy (zarkon):

close

OpenStudy (zarkon):

\[\Large \int\limits_{-1}^{0}(2x^4+8x)^3(8x^3+8)dx=\int\limits_{u(-1)}^{u(0)}u^3\frac{du}{2}\]

OpenStudy (zarkon):

-1 and 0 are limits for x not u

OpenStudy (anonymous):

What's the difference between what you did and I did?

OpenStudy (anonymous):

My calc teacher didn't seem to do that step, kept it the same.

OpenStudy (zarkon):

then he/she is wrong :)

OpenStudy (zarkon):

you should double check your notes. I doubt that they kept them the same...unless they did a work around. But what you wrote is definitely wrong.

OpenStudy (anonymous):

But we will be returning to x no?

OpenStudy (zarkon):

only if you want to

OpenStudy (zarkon):

you can write it in terms of just u or you can return back to x and use the original limits

OpenStudy (anonymous):

Yes, that's the plan. How would I proceed with that?

OpenStudy (zarkon):

find the antiderivative of \(\dfrac{u^3}{2}\)

OpenStudy (anonymous):

\[\frac{ u^4 }{ 8 }\]

OpenStudy (zarkon):

yes. Now replace u by what you set it equal to

OpenStudy (anonymous):

I got -162 for final answer.

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

Can I ask a quick question?

OpenStudy (anonymous):

On something different?

OpenStudy (zarkon):

what?

OpenStudy (anonymous):

To find whether something concaves up or down you'd check the second derivative right?

OpenStudy (zarkon):

if the second derivative exists...yes

OpenStudy (anonymous):

Ok so what do I check it for? I have the second derivative but not sure how to check if it concaves up or down

OpenStudy (zarkon):

it depends on where you evaluate it. it can be concave down on some parts and concave up on others

OpenStudy (anonymous):

Should I make another post?

OpenStudy (zarkon):

prob

OpenStudy (anonymous):

Just made it and tagged you. Thanks